Suppose that the credit remaining on a phone card (in dollars) is a linear function of the total calling time (in minutes). When graphed, the function gives a line with a slope of -0.16
There is $29.00 in credit remaining on the card after 25 minutes of calls. How much credit was there after 18 minutes of calls?

Answer by stanbon(48575)

Suppose that the credit remaining on a phone card (in dollars) is a linear function of the total calling time (in minutes). When graphed, the function gives a line with a slope of -0.16
There is $29.00 in credit remaining on the card after 25 minutes of calls. How much credit was there after 18 minutes of calls?
-----------------
Credit Equation: C(x) = -0.16x + r where "x" is number of minutes of calls
29 = -0.16*25 + r
r = 29+4 dollars
r = 33 dollars
-----------------------
So the Credit Remaining equation is:
C(x) = -0.16x + 33
------------------------
How much credit was there after 18 minutes of calls?
C(18) = -0.16*18+33
C(18) = $30.12
====================
Cheers,
Stan H.